xx CarpetProlongateTest

CarpetProlongateTest

Erik Schnetter <schnetter@cct.lsu.edu>

July 10 2010

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11interpxnx 10Scope: privatex01 INT

11Description: number of points for interpolated region

11  Range  10  x01  Default: 4

110:*  10

11

11interpxny  10Scope: privatex01  INT

11Description: number of points for interpolated region

11  Range  10  x01  Default: 4

110:*  10

11

11interpxnz  10Scope: privatex01  INT

11Description: number of points for interpolated region

11  Range  10  x01  Default: 4

110:*  10

11

11interpxxmax  10Scope: privatex01  REAL

11Description: xmax for interpolated region

11  Range  10  x01  Default: +0.5

11*:*  10

11

11interpxxmin  10Scope: privatex01  REAL

11Description: xmin for interpolated region

11  Range  10  x01  Default: -0.5

11*:*  10

11

11interpxymax  10Scope: privatex01  REAL

11Description: ymax for interpolated region

11  Range  10  x01  Default: +0.5

11*:*  10

11

11interpxymin  10Scope: privatex01  REAL

11Description: ymin for interpolated region

11  Range  10  x01  Default: -0.5

11*:*  10

11

11interpxzmax  10Scope: privatex01  REAL

11Description: zmax for interpolated region

11  Range  10  x01  Default: +0.5

11*:*  10

11

11interpxzmin  10Scope: privatex01  REAL

11Description: zmin for interpolated region

11  Range  10  x01  Default: -0.5

11*:*  10

11

11interpolator  10Scope: privatex01  STRING

11Description: The interpolator to use

11 Range 10 x01 Default: Lagrange polynomial interpolation
11                 10must be a registered interpolator

11interpolatorxoptions 10Scope: privatex01  STRING

11Description: Options for the interpolator

11 Range 10 x01 Default: order=2
11                 10must be a valid options specification

11

11powerxt  10Scope: privatex01  INT

11Description: Polynomial power of t coordinate

11  Range  10  x01  Default: (none)

11*:*  10

11

11powerxx  10Scope: privatex01  INT

11Description: Polynomial power of x coordinate

11  Range  10  x01  Default: (none)

11*:*  10

11

11powerxy  10Scope: privatex01  INT

11Description: Polynomial power of y coordinate

11  Range  10  x01  Default: (none)

11*:*  10

11

11powerxz  10Scope: privatex01  INT

11Description: Polynomial power of z coordinate

11  Range  10  x01  Default: (none)

11*:*  10

11

11prolongation  10Scope: privatex01  KEYWORD

11Description: The prolongation operator to use

11 Range 10 x01 Default: Lagrange
11Lagrange  10

11ENO  10
11WENO  10
11TVD  10

11Lagrangexmonotone

                   10
11

7 Interfaces

General Implements: carpetprolongatetest

Inherits: grid Grid Variables 7.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11  scalar  10  x01  compact110
11                 10  u  x01 description11Grid function
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11 10 x01 tags11ProlongationParameter=”CarpetProlongateTest::prolongation”
11                 10  x01 timelevels113
11                 10  x01variable type11REAL
11  scaled  10  x01  compact110
11                 10  uscaled  x01 description11Scaled grid function
11                 10  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels113
11                 10  x01variable type11REAL
11  difference  10  x01  compact110
11                 10  u0  x01 description11Error in grid function
11                 10  du  x01 dimensions113
11                 10  x01 distribution11DEFAULT
11                 10  x01 group type11GF
11                 10  x01 timelevels113
11                 10  x01variable type11REAL
11 interpxdifference 10  x01  compact110
11                 10  interpxx  x01 description11Error in interpolated grid array
11                 10  interpxy  x01 dimensions113
11                 10  interpxz  x01 distribution11DEFAULT
11                 10  interpxu  x01 group type11ARRAY
11                 10  interpxu0  x01  size11INTERPxNX
11                 10  x  x01  size11INTERPxNY
11                 10  interpxu0  x01  size11INTERPxNZ
11                 10  interpxdu  x01 timelevels111
11                 10  x01variable type11REAL
11 errornorm 10 errornorm x01 compact110
11                 10  x01 description11Norm of error in grid function
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11REAL
11 interpxerrornorm 10 interpxerrornorm x01  compact110
11                 10  x01 description11Norm of error in interpolated grid array
11                 10  x01 dimensions110
11                 10  x01 distribution11CONSTANT
11                 10  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01variable type11REAL

8 Schedule This section lists all the variables which are assigned storage by thorn Carpet/CarpetProlongateTest. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11scalar[3] scaled[3] difference[3]11x
11interpxdifference  11x
11errornorm interpxerrornorm 11x

11x  11x

11

Scheduled Functions

CCTKxINITIAL x carpetprolongatetestxinit x set up initial data x

11x11Language:11fortran
11 x 11Sync:  11scalar
11 x 11x  11scaled
11 x 11Type: 11function
11 x 11Writes: 11u(interior)
11 x 11x  11uscaled(interior)

11

CCTKxEVOL x carpetprolongatetestxinit x set up initial data x

11x11Language:11fortran
11 x 11Reads: 11grid::x
11 x 11x  11grid::y
11 x 11x  11grid::z
11 x 11Sync:  11scalar
11 x 11x  11scaled
11 x 11Type: 11function
11 x 11Writes: 11u(interior)
11 x 11x  11uscaled(interior)

11

MoLxPostStep x carpetprolongatetestxdiff x test data

11x11Language:11fortran
11 x 11Reads: 11grid::x
11 x 11x  11grid::y
11 x 11x  11grid::z

11 x 11x  11u

11 x 11Type: 11function
11 x 11Writes: 11u0(everywhere)
11 x 11x  11du(everywhere)

11

CCTKxINITIAL x carpetprolongatetestxinterpinit x set up interpolation x

11x11Language:11fortran
11 x 11Options: 11global-late
11 x 11Type: 11function
11 x 11Writes: 11interpxx(everywhere)
11 x 11x  11interpxy(everywhere)
11 x 11x  11interpxz(everywhere)

11

MoLxPostStep x carpetprolongatetestxinterp x interpolate x

11x11Language:11fortran
11 x 11Options: 11global-late
11 x 11Reads: 11interpxx
11 x 11x  11interpxy
11 x 11x  11interpxz
11 x 11Type: 11function
11 x 11Writes: 11interpxu(everywhere)

11

MoLxPostStep x carpetprolongatetestxinterpdiff x test interpolated data x

11x11After: 11carpetprolongatetestxinterp
11 x 11Language:11fortran
11 x 11Options: 11global-late
11 x 11Reads: 11interpxx
11 x 11x  11interpxy
11 x 11x  11interpxz
11 x 11x  11interpxu
11 x 11Type: 11function
11 x 11Writes: 11interpxu0(everywhere)
11 x 11x  11interpxdu(everywhere)

11

MoLxPostStep x carpetprolongatetestxnorminit x calculate error norm

11x11After: 11carpetprolongatetestxdiff
11 x 11x  11carpetprolongatetestxinterpdiff
11 x 11Language:11fortran
11 x 11Options: 11global-late
11 x 11Type: 11function
11 x 11Writes: 11errornorm(everywhere)
11 x 11x  11interpxerrornorm(everywhere)

11

MoLxPostStep x carpetprolongatetestxnormcalc x calculate error norm x

11x11After: 11carpetprolongatetestxnorminit
11 x 11Language:11fortran
11 x 11Options: 11global-late
11 x 11x  11loop-local
11 x 11Reads: 11errornorm
11 x 11x  11interpxerrornorm

11 x 11x  11du

11 x 11x  11interpxdu
11 x 11Type: 11function
11 x 11Writes: 11errornorm(everywhere)
11 x 11x  11interpxerrornorm(everywhere)

11

MoLxPostStep x carpetprolongatetestxnormreduce x calculate error norm x

11x11After: 11carpetprolongatetestxnormcalc
11 x 11Language:11fortran
11 x 11Options: 11global-late
11 x 11Reads: 11errornorm
11 x 11x  11interpxerrornorm
11 x 11Type: 11function
11 x 11Writes: 11errornorm(everywhere)
11 x 11x  11interpxerrornorm(everywhere)